OpenStudy (anonymous):

how do you solve (x^3*y)^2x^3/(x^2)^2(y^3)^2

OpenStudy (anonymous):

is it in this form?

OpenStudy (anonymous):

this plus the answer

OpenStudy (anonymous):

\[\frac{\left(x^3 y\right)^2 x^3 \left(y^3\right)^2}{\left(x^2\right)^2}=x^5 y^8 \]